An investment project has an installed cost of $518,297. The cash flows over the 4-year life of the investment are projected to be $287,636, $203,496, $103,802, and $92,556, respectively. What is the NPV of this project if the discount rate is 12.5 percent? Should this project be accepted?
Answer #1

Solution-

Installation cost (Present value of cash flows at 0 period ) = $5,18,297*1= $5,18,297

Present value of cash flows at 1 period = $287,636/1.125^1= $255,676.44

Present value of cash flows at 2 period = $203,496/1.125^2= $160,786.96

Present value of cash flows at 3 period = $103,802/1.125^3= $72,903.46

Present value of cash flows at 4 period = $92,556/1.125^4= $57,782.25

NPV = Present value of cash inflows - Present value of cash outflow

NPV = $255,676.44 + $160,786.96 + $72,903.46 + $57,782.25 - $518,297= $28,852.11

Project should be accepted as there is positive NPV.